OpenStudy (anonymous):

how to find the stationary point of the curve when y=xe^-2x

OpenStudy (anonymous):

sorry i had to repost the question because my earlier question was stated wrong :D

OpenStudy (anonymous):

Derivative e^-2x(1-2x) -> x= 1/2 for max

OpenStudy (anonymous):

i found out that derivative is -2e^-2x is that right? if so where do i go from here to get x=1/2

OpenStudy (anonymous):

i found this (x x)/e^2 = 0 at x = 0

OpenStudy (anonymous):

and thats for minimum btw

OpenStudy (anonymous):

The derivative is what I said, which gives you the half.

OpenStudy (anonymous):

I think its important to know Stationary points of a curve ar the points where, for example, if you move from left to right,the curve is increasing, reaches a maximum point, stops, and then starts decreasing, the very top point (the maximum) is the stationary point, there can also be the minimimum point (if the curve is like a valley). this equates to a slope of zero, hence, reason for finding derivative here, which gives you slope.

OpenStudy (anonymous):

You can find the concavity with a second derivative (derivative of the derivative).